Exercise 5.71 R The zine within a copper-plated penny will dissolve in hydrochloric acid if the copper coating is filed down circle answers
Answer #1

According to ideal gas equation, PV= m RT. 02 P Pressure V Volume m Mass My Molar mass R Gas Constant I absolute temperature.Molar mass of He → 2.0169/mol T= 25°C - 6 V=0.9572 P=0.986 bar = 25+273K I bar = 0.987 = 298K atm P=0.986X0.987 = 0.973 atom